OpenStudy (anonymous):

given ln(10)=2.3, use differentials to approximate ln(10.2)

OpenStudy (anonymous):

This is the same procedure as with using a linear approximation. To approximate a function at some value \(x\) near a known \(a\), you have \[\begin{align*} f(x)&\approx f'(a)(x-a)+f(a) \end{align*}\] \(f(x)=\ln x~~\implies ~~f'(x)=\dfrac{1}{x}\). In this case, \(a=10\) and \(x=10.2\). \[\begin{align*} \ln10.2&\approx \frac{1}{10}(10.2-10)+2.3\\ &=0.02+2.3\\ &=2.32 \end{align*}\]